Blip operations 02 01 11
Download
1 / 3

BLIP operations 02-01-11 - PowerPoint PPT Presentation


  • 81 Views
  • Uploaded on

BLIP operations 02-01-11. Status – running at 117.5MeV BLIP weekly average beam increased to 96.1µA, highest this year. Reliability very good with total operating time equal to 162.9 hours. Operations issues Investigation of cause of Ga targets leaks continues.

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' BLIP operations 02-01-11' - kyrene


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
Blip operations 02 01 11
BLIP operations 02-01-11

Status – running at 117.5MeV

BLIP weekly average beam increased to 96.1µA, highest this year. Reliability very good with total operating time equal to 162.9 hours.

Operations issues

Investigation of cause of Ga targets leaks continues.

No Ga targets in beam now. Replaced with an Al disk.

There is indirect evidence from our recent research irradiations that the beam spot is smaller this year, thus increasing the power density. We performed a high resolution phosphor image plate measurement of the beam profile that showed this year’s beam spot was 4% narrower than last year. It is not clear if this is the smoking gun . The same Ga targets have been successfully irradiated at LANL with 2.5 times more beam intensity than we have but with a raster beam creating a more uniform intensity profile. So we are pursuing a broader Gausssian tune at this time.

Linac prepared a new tune that was measured to actually be narrower than before. We are trying again today.

We will shutdown tomorrow for ~3.5 hours to collect 400mL of cooling water from which we will chemically extract Be-7 for ORNL research.


Leaking ga target prepared by lanl
Leaking Ga target prepared by LANL

Burn marks on BNL fabricated targets are similar, and always off center.


ad